# The Most Common Habitable Planets III – Modeling Temperature Forcing and Surface Conditions on Rocky Exoplanets and Exomoons

### v1

@@ 4.1 Selection criteria
We have selected two exoplanets from all the confirmed exoplanets available at the NASA Exoplanet Archive Database [^2] https://exoplanetarchive.ipac.caltech.edu/ that we classified as potentially rocky and are inside the classical HZ of their stars, with the assumptions used in 36 and 37. We calculated the planetary density for the subgroup of exoplanets that present measured mass values using the results from the interior planetary models [^3] https://lweb.cfa.harvard.edu/ lzeng/planetmodels.html 86; 87. We considered as potentially rocky planets those that meet both of the following conditions: planetary density between 4 g/cm3 and 10 g/cm3 and planetary radius less than $2\,R_{\oplus}$.
The HZ around the stars were estimated according to the formalism presented in Kopparapu et al. 37, although other definitions exist in the literature as, for example, in Ramirez & Kaltenegger 64 and Pierrehumbert & Gaidos 58. The effective flux received by the planet (Seff) was calculated considering the effective temperature of the star (Teff) and the coefficients $a$, $b$, $c$ and $d$, taken from Kopparapu et al. 38 when considering an exoplanet with $1\,M_{\oplus}$. The limits for the inner and outer borders of HZ calculation were chosen as the runaway greenhouse and the maximum greenhouse effect, respectively.
@@
We have used a fixed value of eccentricity of $e=0.1$ in our calculations, which is within the uncertainties for the measured minimum eccentricities ($0.04^{+0.07}_{-0.04}$ and $0.04^{+0.08}_{-0.04}$ for Kepler-186 f and Kepler-442 b, respectively 78) and a sizable enough value to produce strong results from our model. We have also simulated different atmospheric thicknesses and albedos, as shown in Table 2. Albedo values were chosen to be in rough agreement with the types of atmosphere under consideration and, therefore, we used as reference Mars’ and Earth’s mean albedos, which are 0.25 and 0.30, respectively 57. We also included the case of albedo $A=0.40$ to take into consideration possible more icy or cloudy exoplanets. We considered an albedo of $0.27$ to model Titan’s atmosphere 53.
@@
Table 1: Radius ($R_{P}$), semi-major axis ($a$), minimum orbital eccentricity ($e_{min}$) and orbital period ($P$) for exoplanets Kepler-186 f and Kepler-442 b and host stars’ mass ($M_{\star}$), radius ($R_{\star}$), effective temperature ($T_{eff}$), luminosity ($L_{\star}$) and metallicity [Fe/H]. Data were taken from Torres et al. 78.
@@ 4.2 Results
Figure 5: Brightness temperature profiles for Kepler-186 f (lower set of curves) and Kepler-442 b (upper set of curves) for each model in Table 2. Dotted brown curves represent model 1, solid blue curves represent model 2, dashed orange curves represent model 3 and dot-dashed pink curves represent model 4.
@@ 5 A first look at exomoons
In the rest of this section, we model moons in circular orbits around their planets, which is a reasonable approach since they are expected to be in tidal lock. We do not take into account the effect of eclipses, when the planet blocks the stellar flux received by the moon, since we consider that these events will last a negligible amount of time, if compared with the orbital periods considered 12. In fact, just like for the Galilean moons of Jupiter, for general orbits aligned at the same plane, eclipses do not affect much the total flux received. We also verified that the center of mass of the planet-moon system lies inside the radius of the planet in all cases, which means that a model of the moons orbiting the center of the planets can be considered as a good approximation. We further neglect the contribution of planetary illumination, since we focus on exomoons within the stellar HZs, for which stellar illumination will strongly dominate over the planetary one, at least for old, evolved systems. Even though this contribution may be relevant in certain cases, it can be reasonably expected not to introduce any serious effect in our results given the scenarios we model see 12; 88.
@@ 5.1 Tests with Solar System’s moons
In order to show the agreement of our results with actual measurements, we applied the procedure described in Section 2 for Titan, Ganymede and Europa. The values used in our model for each moon are shown in Table 3. We assumed Ganymede and Europa have no atmosphere and Titan has an atmosphere with $\chi=5\times 10^{6}$ mol/m2 and $c_{p}=32$ mol-1 K-1, as already described in Section 3. All moons were assumed to have a circular orbit around the planet and we used for Saturn a mass of $95.16\,M_{\oplus}$, a semi-major axis of $9.583$ AU and an eccentricity of $0.0565$. For Jupiter, we used adopted a mass of $M_{J}=317.83\,M_{\oplus}$ a semi-major axis of $5.204$ AU and an eccentricity of $0.0489$. All these values we taken from NASA’s “fact sheet" webpage.
Table 3: Mass, semi-major axis, Bond albedo and assumed atmospheric column density for moons in the Solar System. A value of $\chi=0$ means we assumed the moon has no atmosphere. Values of mass and semi-major axis were taken from the NASA’s “fact sheet" webpage. References for the values of Bond albedo are Neff et al. 53 for Titan, Squyres & Veverka 75 for Ganymede and Buratti & Veverka 6 for Europa.
@@
The results can be seen in Fig. 6, in which higher frequency oscillations due to the moons’ orbits around the planets are superimposed to lower frequency ones, due to the planets’ orbits around the star. The brightness temperature ranges predicted for each moon can be compared with the measured values of the surface temperatures for each one: 89 K $<T<$ 94 K for Titan 34, 75 K $<T<$ 152 K for Ganymede 74; 55 and 86 K $<T<$ 132 K for Europa 73. Notice that, the brightness and surface temperatures should be the same only for the case of no atmosphere, which may be a good approximation for Europa and Ganymede, but does not apply to Titan. Also, Europa, with a rotation period of 3.53.5~{}3.53.5 days, is the only moon that can be considered as a fast rotator among the three. Nevertheless, our model produces temperature ranges similar to measurements for these three moons. This lends confidence that our model may realistically represent cases for which conditions better approach our assumptions.
@@ 5.2 The case of circular planetary orbits
This choice of parameters likely represents extremes, and should cover a wide range of real cases. Jupiter-sized planets are scarce around red dwarfs see, for example, 40, and a Ganymede-sized satellite is probably as large an exomoon as can be expected. In fact, it seems that more massive and larger planets are likelier to form around more massive stars 44 and also to host larger and more massive moons 47; 15. A Mars-sized moon (four times as massive as Ganymede) is a reasonable assumption 77, though such an object will not necessarily resemble the composition of Mars, and can be expected to have a more mixed rock/ice makeup.
@@
The radii of the planet’s orbits are chosen so that they receive from the star the same flux received by Earth from the Sun. They are 0.14 AU and 1.44 AU, for cases (i) and (ii), respectively. The radii of the moon’s orbits are chosen in order that they have a period of 3 days around each planet, so they can be considered as fast rotators. This results in radii of 0.004 AU and 0.006 AU, for cases (i) and (ii) respectively. We have determined the Roche radius for these hypothetical planets and exomoons assuming that the moon with the mass of Ganymede has also Ganymede’s density and the moon with the mass of Mars has Mars’ density. The results are $0.001$ AU for both cases, so we can be assured that their orbital radius lie beyond the Roche limit. The Hill radii are estimated to be 0.009 AU and 0.01 AU, for cases (i) and (ii), respectively, so the satellites lie within the stability zones of their planets 16.
@@
Figure 7: Brightness temperature profile for the hypothetical exomoons studied in Section 5.2. Blue curves are obtained for the lower mass moon (Ganymede’s mass), and red curves are obtained for the more massive moon (Mars’ mass). Each panel presents the results for the type of atmosphere and albedo indicated.
@@ 5.3 Eccentric planetary orbits
Let us now study the case of a hypothetical exomoon in circular orbit around a planet with eccentric orbit around its star. We studied a moon with the mass of Mars ($M_{{\rm exomoon}}=0.107\,M_{\oplus}$) in a circular orbit around a planet with mass of three times that of Jupiter, which was the case found to yield the most conservative results (smallest temperature variations) in the previous section. The radius of the moon’s orbit is determined in order for its period to be of 3 days. This results in an orbital radius of $0.006$ AU, which is $\approx 0.8$ times that of Ganymede’s semi-major axis, for example. The Roche limit for this case is 0.001 AU. In this section, we studied the case of four different stellar types, adding K and G dwarfs to the M and F dwarfs studied in the previous section. This is because planets orbiting M dwarfs are expected to be in circular orbits and in this section we want to study the effects of planetary eccentric orbits. We selected the planet’s semi-major axis so that the flux it would receive from the star at the apsis would be the same as the flux Earth receives from the Sun. Table 4 shows the characteristics of the selected stars that we chose to represent each stellar type and the value of the semi-major axis obtained. The values quoted for Gliese 687 and HD 219134 were taken from Pinotti 59 and the values for HD 115383 were taken from Da Silva et al. 9. The Sun’s effective temperature was taken from Prša et al. 61.
@@
For the F and G dwarfs, we applied our methodology for three values of planetary eccentricity, $e_{P}$ = 0, 0.3 and 0.6. For the M dwarf, we considered only a planet in circular orbit and for the K dwarf, we adopted eccentricities of 0 and 0.3. This choice was made because, specially for M dwarfs but also for K dwarfs, most orbits inside the HZ are expected to have been circularized and, also, the fact that the HZ is narrower and closer to the star makes planets with high eccentricities less interesting from an astrobiological point of view. The Hill radius for these cases ranges from 0.018 AU to 0.133 AU, and the value of the radius for our hypothetical exomoons is within the stability zone in all cases studied 16. We studied the case of no atmosphere and albedo 0.25 and an Earth like atmosphere, described in Table 2 as model 2, with $\chi=3\times 10^{4}$ mol/m2 and albedo 0.3. In fact, although very thin atmospheres should be expected at those distances from the host star, we consider the possibility that the exomoons might have kept a somewhat thick atmosphere acquired at larger distances from the star, before planetary migration, if it is able to replenish its volatile inventory by continued volcanic activity, tidally induced.
@@
Figure 8: Brightness temperature profile for the hypothetical exomoons, with no atmosphere and albedo 0.25, studied in Section 5.3. Different colors represent moons in different planetary systems with an M dwarf (pink), K dwarf (brown), F dwarf (orange) and G dwarf (blue). The upper panel shows the case for a planet in circular orbit ($e_{p}=0$), the middle panel shows the same for a planetary eccentricity of $e_{p}$ = 0.3 and the lower panel shows the same for a planetary eccentricity of $e_{p}$ = 0.6.
@@ 5.4 Hypothetical exomoons around real exoplanets
Uncertainties of the measured stellar and planetary parameters (with the exception of eccentricity) are mostly below $3\%$ and eccentricity measurements are the dominant source of error in the curves shown in Figs. 10 and 11. The luminosity of HD 165155, $0.70\pm 0.08\,L_{\odot}$, is an exception with a $\sim 10\%$ error bar 33. Considering these uncertainties, the lower temperature variation of the hypothetical exomoon could reach $\sim 10$ K, for HD 221287 b’s case, and the higher $\sim 160$ K for HD 128356 b’s case.
@@ 6 Tidal dissipation in massive habitable exomoons
The potential habitability of exomoons has long become an exciting topic in planetary science and astrobiology 15; 42; 49; 12; 88, the chief reason being the expectation that massive exomoons are likely to exist around massive gas giants possibly up to one-third of an Earth mass 15. These large bodies could maintain liquid water at their surface for periods commensurate with the evolution of microbial life and, moreover, 14 report that, for orbital periods compatible with the HZ of FGK stars, exomoons around giant planets are very likely to remain in stable orbits. Due to observational biases the first exomoons to be detected will be large, possibly more massive than Mars 25.
Current evidence suggests that Mars itself was probably habitable for longer than 1 Gyr though perhaps only episodically, 62; 85 and that volcanism was a relevant contributing factor 65; 84. The maintenance of a geophysical internal engine is, probably, as relevant to providing long-lived habitable conditions on the surface of rocky exomoons as to supporting subsurface oceans or liquid habitats in icy ones. Subsurface biospheres may be longer lasting and perhaps even more benign, but such biospheres will be much harder to detect on exomoons than those based on surface conditions, which are the ones of interest to our approach.
Continued dissipation of tidal energy in large, rocky exomoons can be expected to maintain expressive volcanic activity, helping sustain a sizable atmosphere and maintain habitability, even in the absence of plate tectonics. Such activity might not be unlike the recently suggested protracted episodes of supervolcanism on Mars 81, which may have lasted a few hundred Myr. Solar System data clearly indicates that there is a minimum required mass for rock bodies to hold on to a substantial fraction of their volatile content over long timescales, and that Mars-sized objects are below this threshold unless positioned very far from outer limit of the HZ, as testified by the case of Titan. The exomoon scenarios we invoke are severely affected as they assume low-mass objects inside the HZ and therefore closer to the host star than most large moons in our Solar System. Therefore, exogenous sources such as tidal dissipation enabling the maintenance of geologic activity and volcanism are probably indispensable in allowing exomoons, within the HZ of FGKM stars, to keep substantial atmospheres over long timescales. Tidal dissipation has been found to be a relevant energy source for Mars-sized exomoons inside the HZ of such stars 12; 88.
Towards evaluating some of the exomoon-exoplanet scenarios we investigated in Sec. 5, in terms of surface energy fluxes from tidal dissipation, compared to observed cases in our Solar System, we estimated the rates of energy dissipated per unit area at the surface of hypothetical exomoons, along one orbital period, from the formalism (their equations 4.77 and 4.197) of 52:
@@
We aim at providing single-body only, order of magnitude estimates in a preliminary exploration, so that the exomoon cases we examined in Sec. 5 can be put into context. We do not consider details of the satellite architecture and interaction with companion exomoons. The values of $\mu$ were retrieved from 52 and correspond to $\mu_{\rm rocky}$ = 5.0 $\times$ 106 dyne.cm-2 for $\rho$ $\sim$ 3 g.cm3. We reproduced the values of $\dot{E}$ given by 52 for the actual cases of Io, Europa and Ganymede to within a factor of two or better, differences probably being accounted for by different assumptions about $Q_{\rm p}$. Our estimates of $F_{\rm tidal}$ assume $Q_{\rm p}$ = 100, a value comparable to the one commonly cited for Mars ($Q_{\rm p}$ = 86) and are listed in Table 7, along with the corresponding values for Io, Europa and Ganymede.

### v2

@@ 4.1 Selection criteria
We have selected two exoplanets from all the confirmed exoplanets available at the NASA Exoplanet Archive Database [^2] https://exoplanetarchive.ipac.caltech.edu/ that we classified as potentially rocky and are inside the classical HZ of their stars, with the assumptions used in 34 and 35. We calculated the planetary density for the subgroup of exoplanets that present measured mass values using the results from the interior planetary models 84; 85. We considered as potentially rocky planets those that meet both of the following conditions: planetary density between 4 g/cm3 and 10 g/cm3 and planetary radius less than $2\,R_{\oplus}$.
The HZ around the stars were estimated according to the formalism presented in Kopparapu et al. 35, although other definitions exist in the literature as, for example, in Ramirez & Kaltenegger 62 and Pierrehumbert & Gaidos 56. The effective flux received by the planet (Seff) was calculated considering the effective temperature of the star (Teff) and the coefficients $a$, $b$, $c$ and $d$, taken from Kopparapu et al. 36 when considering an exoplanet with $1\,M_{\oplus}$. The limits for the inner and outer borders of HZ calculation were chosen as the runaway greenhouse and the maximum greenhouse effect, respectively.
@@
We have used a fixed value of eccentricity of $e=0.1$ in our calculations, which is within the uncertainties for the measured minimum eccentricities ($0.04^{+0.07}_{-0.04}$ and $0.04^{+0.08}_{-0.04}$ for Kepler-186 f and Kepler-442 b, respectively 76) and a sizable enough value to produce strong results from our model. We have also simulated different atmospheric thicknesses and albedos, as shown in Table 2. Albedo values were chosen to be in rough agreement with the types of atmosphere under consideration and, therefore, we used as reference Mars’ and Earth’s mean albedos, which are 0.25 and 0.30, respectively 55. We also included the case of albedo $A=0.40$ to take into consideration possible more icy or cloudy exoplanets. We considered an albedo of $0.27$ to model Titan’s atmosphere 51.
@@
Table 1: Radius ($R_{P}$), semi-major axis ($a$), minimum orbital eccentricity ($e_{min}$) and orbital period ($P$) for exoplanets Kepler-186 f and Kepler-442 b and host stars’ mass ($M_{\star}$), radius ($R_{\star}$), effective temperature ($T_{eff}$), luminosity ($L_{\star}$) and metallicity [Fe/H]. Data were taken from Torres et al. 76.
@@ 4.2 Results
Figure 5: Brightness temperature profiles for Kepler-186 f (lower set of curves) and Kepler-442 b (upper set of curves) for each model in Table 2. Dotted brown curves represent model 1, solid curves represent model 2, dashed curves represent model 3 and dot-dashed curves represent model 4.
@@ 5 A first look at exomoons
In the rest of this section, we model moons in circular orbits around their planets, which is a reasonable approach since they are expected to be in tidal lock. We do not take into account the effect of eclipses, when the planet blocks the stellar flux received by the moon, since we consider that these events will last a negligible amount of time, if compared with the orbital periods considered 12. In fact, just like for the Galilean moons of Jupiter, for general orbits aligned at the same plane, eclipses do not affect much the total flux received. We also verified that the center of mass of the planet-moon system lies inside the radius of the planet in all cases, which means that a model of the moons orbiting the center of the planets can be considered as a good approximation. We further neglect the contribution of planetary illumination, since we focus on exomoons within the stellar HZs, for which stellar illumination will strongly dominate over the planetary one, at least for old, evolved systems. Even though this contribution may be relevant in certain cases, it can be reasonably expected not to introduce any serious effect in our results given the scenarios we model see 12; 86.
@@ 5.1 Tests with Solar System’s moons
In order to show the agreement of our results with actual measurements, we applied the procedure described in Section 2 for Titan, Ganymede and Europa. The values used in our model for each moon are shown in Table 3. We assumed Ganymede and Europa have no atmosphere and Titan has an atmosphere with $\chi=5\times 10^{6}$ mol/m2 and $c_{p}=32$ mol-1 K-1, as already described in Section 3. All moons were assumed to have a circular orbit around the planet and we used for Saturn a mass of $95.16\,M_{\oplus}$, a semi-major axis of $9.583$ AU and an eccentricity of $0.0565$. For Jupiter, we used adopted a mass of $M_{J}=317.83\,M_{\oplus}$ a semi-major axis of $5.204$ AU and an eccentricity of $0.0489$. All these values we taken from NASA’s “planetary fact sheet" webpage.
Table 3: Mass, semi-major axis, Bond albedo and assumed atmospheric column density for moons in the Solar System. A value of $\chi=0$ means we assumed the moon has no atmosphere. Values of mass and semi-major axis were taken from the NASA’s “planetary fact sheet" webpage. References for the values of Bond albedo are Neff et al. 51 for Titan, Squyres & Veverka 73 for Ganymede and Buratti & Veverka 6 for Europa.
@@
The results can be seen in Fig. 6, in which higher frequency oscillations due to the moons’ orbits around the planets are superimposed to lower frequency ones, due to the planets’ orbits around the star. The brightness temperature ranges predicted for each moon can be compared with the measured values of the surface temperatures for each one: 89 K $<T<$ 94 K for Titan 33, 75 K $<T<$ 152 K for Ganymede 72; 53 and 86 K $<T<$ 132 K for Europa 71. Notice that, the brightness and surface temperatures should be the same only for the case of no atmosphere, which may be a good approximation for Europa and Ganymede, but does not apply to Titan. Also, Europa, with a rotation period of 3.53.5~{}3.53.5 days, is the only moon that can be considered as a fast rotator among the three. Nevertheless, our model produces temperature ranges similar to measurements for these three moons. This lends confidence that our model may realistically represent cases for which conditions better approach our assumptions.
@@ 5.2 The case of circular planetary orbits
This choice of parameters likely represents extremes, and should cover a wide range of real cases. Jupiter-sized planets are scarce around red dwarfs see, for example, 38, and a Ganymede-sized satellite is probably as large an exomoon as can be expected. In fact, it seems that more massive and larger planets are likelier to form around more massive stars 42 and also to host larger and more massive moons 45; 14. A Mars-sized moon (four times as massive as Ganymede) is a reasonable assumption 75, though such an object will not necessarily resemble the composition of Mars, and can be expected to have a more mixed rock/ice makeup.
@@
The radii of the planet’s orbits are chosen so that they receive from the star the same flux received by Earth from the Sun. They are 0.14 AU and 1.44 AU, for cases (i) and (ii), respectively. The radii of the moon’s orbits are chosen in order that they have a period of 3 days around each planet, so they can be considered as fast rotators. This results in radii of 0.004 AU and 0.006 AU, for cases (i) and (ii) respectively. We have determined the Roche radius for these hypothetical planets and exomoons assuming that the moon with the mass of Ganymede has also Ganymede’s density and the moon with the mass of Mars has Mars’ density. The results are $0.001$ AU for both cases, so we can be assured that their orbital radius lie beyond the Roche limit. The Hill radii are estimated to be 0.009 AU and 0.01 AU, for cases (i) and (ii), respectively, so the satellites lie within the stability zones of their planets 15.
@@
Figure 7: Brightness temperature profile for the hypothetical exomoons studied in Section 5.2. For each panel, curves with the highest amplitudes are obtained for the lower mass moon (Ganymede’s mass), and curves with the lowest ones are obtained for the more massive moon (Mars’ mass). Each panel presents the results for the type of atmosphere and albedo indicated.
@@ 5.3 Eccentric planetary orbits
Let us now study the case of a hypothetical exomoon in circular orbit around a planet with eccentric orbit around its star. We studied a moon with the mass of Mars ($M_{{\rm exomoon}}=0.107\,M_{\oplus}$) in a circular orbit around a planet with mass of three times that of Jupiter, which was the case found to yield the most conservative results (smallest temperature variations) in the previous section. The radius of the moon’s orbit is determined in order for its period to be of 3 days. This results in an orbital radius of $0.006$ AU, which is $\approx 0.8$ times that of Ganymede’s semi-major axis, for example. The Roche limit for this case is 0.001 AU. In this section, we studied the case of four different stellar types, adding K and G dwarfs to the M and F dwarfs studied in the previous section. This is because planets orbiting M dwarfs are expected to be in circular orbits and in this section we want to study the effects of planetary eccentric orbits. We selected the planet’s semi-major axis so that the flux it would receive from the star at the apsis would be the same as the flux Earth receives from the Sun. Table 4 shows the characteristics of the selected stars that we chose to represent each stellar type and the value of the semi-major axis obtained. The values quoted for Gliese 687 and HD 219134 were taken from Pinotti 57 and the values for HD 115383 were taken from Da Silva et al. 9. The Sun’s effective temperature was taken from Prša et al. 59.
@@
For the F and G dwarfs, we applied our methodology for three values of planetary eccentricity, $e_{P}$ = 0, 0.3 and 0.6. For the M dwarf, we considered only a planet in circular orbit and for the K dwarf, we adopted eccentricities of 0 and 0.3. This choice was made because, specially for M dwarfs but also for K dwarfs, most orbits inside the HZ are expected to have been circularized and, also, the fact that the HZ is narrower and closer to the star makes planets with high eccentricities less interesting from an astrobiological point of view. The Hill radius for these cases ranges from 0.018 AU to 0.133 AU, and the value of the radius for our hypothetical exomoons is within the stability zone in all cases studied 15. We studied the case of no atmosphere and albedo 0.25 and an Earth like atmosphere, described in Table 2 as model 2, with $\chi=3\times 10^{4}$ mol/m2 and albedo 0.3. In fact, although very thin atmospheres should be expected at those distances from the host star, we consider the possibility that the exomoons might have kept a somewhat thick atmosphere acquired at larger distances from the star, before planetary migration, if it is able to replenish its volatile inventory by continued volcanic activity, tidally induced.
@@
Refer to captionRefer to captionRefer to caption
Figure 8: Brightness temperature profile for the hypothetical exomoons, with no atmosphere and albedo 0.25, studied in Section 5.3. Different colors represent moons in different planetary systems with an M dwarf, K dwarf, F dwarf and G dwarf. The upper panel shows the case for a planet in circular orbit ($e_{p}=0$), the middle panel shows the same for a planetary eccentricity of $e_{p}$ = 0.3 and the lower panel shows the same for a planetary eccentricity of $e_{p}$ = 0.6.
Refer to captionRefer to captionRefer to caption
@@ 5.4 Hypothetical exomoons around real exoplanets
Uncertainties of the measured stellar and planetary parameters (with the exception of eccentricity) are mostly below $3\%$ and eccentricity measurements are the dominant source of error in the curves shown in Figs. 10 and 11. The luminosity of HD 165155, $0.70\pm 0.08\,L_{\odot}$, is an exception with a $\sim 10\%$ error bar 32. Considering these uncertainties, the lower temperature variation of the hypothetical exomoon could reach $\sim 10$ K, for HD 221287 b’s case, and the higher $\sim 160$ K for HD 128356 b’s case.
@@ 6 Tidal dissipation in massive habitable exomoons
The potential habitability of exomoons has long become an exciting topic in planetary science and astrobiology 14; 40; 47; 12; 86, the chief reason being the expectation that massive exomoons are likely to exist around massive gas giants possibly up to one-third of an Earth mass 14. These large bodies could maintain liquid water at their surface for periods commensurate with the evolution of microbial life and, moreover, 13 report that, for orbital periods compatible with the HZ of FGK stars, exomoons around giant planets are very likely to remain in stable orbits. Due to observational biases the first exomoons to be detected will be large, possibly more massive than Mars 24.
Current evidence suggests that Mars itself was probably habitable for longer than 1 Gyr though perhaps only episodically, 60; 83 and that volcanism was a relevant contributing factor 63; 82. The maintenance of a geophysical internal engine is, probably, as relevant to providing long-lived habitable conditions on the surface of rocky exomoons as to supporting subsurface oceans or liquid habitats in icy ones. Subsurface biospheres may be longer lasting and perhaps even more benign, but such biospheres will be much harder to detect on exomoons than those based on surface conditions, which are the ones of interest to our approach.
Continued dissipation of tidal energy in large, rocky exomoons can be expected to maintain expressive volcanic activity, helping sustain a sizable atmosphere and maintain habitability, even in the absence of plate tectonics. Such activity might not be unlike the recently suggested protracted episodes of supervolcanism on Mars 79, which may have lasted a few hundred Myr. Solar System data clearly indicates that there is a minimum required mass for rock bodies to hold on to a substantial fraction of their volatile content over long timescales, and that Mars-sized objects are below this threshold unless positioned very far from outer limit of the HZ, as testified by the case of Titan. The exomoon scenarios we invoke are severely affected as they assume low-mass objects inside the HZ and therefore closer to the host star than most large moons in our Solar System. Therefore, exogenous sources such as tidal dissipation enabling the maintenance of geologic activity and volcanism are probably indispensable in allowing exomoons, within the HZ of FGKM stars, to keep substantial atmospheres over long timescales. Tidal dissipation has been found to be a relevant energy source for Mars-sized exomoons inside the HZ of such stars 12; 86.
Towards evaluating some of the exomoon-exoplanet scenarios we investigated in Sec. 5, in terms of surface energy fluxes from tidal dissipation, compared to observed cases in our Solar System, we estimated the rates of energy dissipated per unit area at the surface of hypothetical exomoons, along one orbital period, from the formalism (their equations 4.77 and 4.197) of 50:
@@
We aim at providing single-body only, order of magnitude estimates in a preliminary exploration, so that the exomoon cases we examined in Sec. 5 can be put into context. We do not consider details of the satellite architecture and interaction with companion exomoons. The values of $\mu$ were retrieved from 50 and correspond to $\mu_{\rm rocky}$ = 5.0 $\times$ 106 dyne.cm-2 for $\rho$ $\sim$ 3 g.cm3. We reproduced the values of $\dot{E}$ given by 50 for the actual cases of Io, Europa and Ganymede to within a factor of two or better, differences probably being accounted for by different assumptions about $Q_{\rm p}$. Our estimates of $F_{\rm tidal}$ assume $Q_{\rm p}$ = 100, a value comparable to the one commonly cited for Mars ($Q_{\rm p}$ = 86) and are listed in Table 7, along with the corresponding values for Io, Europa and Ganymede.